# Uncertainty-aware Evidential Fusion-based Learning for Semi-supervised Medical Image Segmentation

### 09 Apr 2024 v1

@@ Abstract.
Figure 1. The diagram illustrates the fusion process of the model’s predictions for the voxels at the same location in the original samples and the samples after mixup restoration. It demonstrates the redistribution and adjustment of the voxel’s confidence level in belonging to a category after interacting with the overall uncertainty of the prediction.
@@
  • The proposed method realizes the state-of-the-art (SOTA) performance on the task of semi-supervised medical image segmentation compared with previous advanced approaches on four popular benchmark datasets.
@@ 2.1. Generalized Probabilistic Framework
Evidential deep learning has achieved excellent results in multiple fields. We extend the framework of evidential deep learning to semi-supervised medical segmentation tasks by further improving the uncertainty measurement methods and probability assignments fusion rule of traditional evidence theory. We map the uncertainty measures of evidential deep learning to the multi-objective subsets of traditional evidence theory and fuse reliable information for both mixed and original sample annotation areas during the prediction phase, thereby achieving more robust model predictions with refined uncertainty measures. Moreover, inspired by curriculum learning (Bengio et al., 2009), we design a novel learning mechanism based on the refined uncertainty measure for the model to gradually focus on difficult-to-learn features.
In detail, let $g^{(x,y,z)}_{\mathcal{M}^{l}},g^{(x,y,z)}_{\mathcal{T}^{l_{1}}},g^{(x,y,z)}_{\mathcal{T}^{l_{2}}},g^{(x,y,z)}_{\mathcal{M}^{ul}},g^{(x,y,z)}_{\mathcal{T}^{u}}\in\mathcal{R}^{C}$ denote belief mass from evidential predictive results (Sensoy et al., 2018; Deng et al., 2023) from classifier of initial network $\mathcal{N}$ in pre-training stage on samples $\mathcal{M}^{l},\mathcal{T}^{l}$, and student network $\mathcal{N}_{s}$ in self-training stage on samples $\mathcal{T}^{l},\mathcal{M}^{ul},\mathcal{T}^{u}$ for voxel at position $(x,y,z)$. The final trained parameters of the initial network $\mathcal{N}$ will be utilized as the parameters of the teacher and student networks in the beginning of the self-training phase. The basic probability mass assignments for voxels can be defined as $\mathcal{P}^{(x,y,z)}=\{\{g^{(x,y,z)}(C_{n})\}_{n=0}^{N-1},u^{(x,y,z)}\}$ and $u^{(x,y,z)}=1-\sum_{n=0}^{N-1}g^{(x,y,z)}(C_{n})$ denotes the original uncertainty measure. The generalized form of predictive probability assignments can be formulated as:
@@ 2.2. Uncertainty-aware Evidential Fusion
In the previous section, we have mentioned that the method proposed in this article is divided into pre-training and self-training stages. For the pre-training stage, we first feed the samples into the initial network to obtain their predictive results. Due to the uniqueness of the mixed strategy, we match the original labeled samples with the mixed samples one-to-one, and merge the evidential results of the corresponding areas of the two predictions. More specifically, if there are samples $E$ and $T$, which exchange features in the corresponding area $\mathcal{D}$ to obtain the mixed samples $E^{\prime}$ and $T^{\prime}$, after obtaining the evidential predictive results for the original samples $E$, $T$ and the mixed samples $E^{\prime}$, $T^{\prime}$, we return the exchanged areas of the mixed samples prediction back to their original positions and fuse them with the model’s predictive results for the original samples. Finally, the combined results and prediction of original data are used for loss calculation with the corresponding labels $Y^{l}$ to obtain $\mathcal{L}^{l_{1}}$, both of which consist of Dice loss and Cross-Entropy loss. We consider the fusion of evidential results from two different predictive outputs that come from the same sample and are located at the same position $(x,y,z)$. This process can be defined as:
@@
(3) $g_{fused}(C_{K})=g_{\mathcal{M}^{l}}(C_{K})g_{\mathcal{T}^{l}}(C_{K})+\frac{|C_{K}|}{|C_{N}|}\cdot$
$(g_{\mathcal{M}^{l}}(C_{K})g_{\mathcal{T}^{l}}(C_{N})+g_{\mathcal{T}^{l}}(C_{K})g_{\mathcal{M}^{l}}(C_{N}))$
where $K=0,...,N-1$, $|C_{K}|=1,|C_{K}|=N$ and $g_{fused}(C_{N})=g_{\mathcal{M}^{l}}(C_{N})*g_{\mathcal{T}^{l}}(C_{N})$. This fusion process evolves from the traditional Dempster’s combination rule requiring normalization. The proposed fusion strategy maintains the design for the interaction between the confidence level of each category and the uncertain information, while simultaneously controlling the degree of interaction to avoid the confidence level of each category being dominated by uncertain information. Therefore, the interaction between the confidence associated with each category and the uncertain components is subject to certain restrictions, and we prefer the model to reduce uncertainty through the learning process, rather than directly converting uncertain parts into definite confidence assignments. If the fusion of probabilities follows the original Dempster’s combination rule directly (Dempster, 2008), it may lead to the model being overly confident about certain predictions, thereby diminishing the focus on corresponding features, even though learning these features should still be encouraged. Similarly, for samples in the self-training stage, the fusion process they undergo is very similar to that in the pre-training stage. Labeled and unlabeled data undergo the same mixing operation as in the pre-training stage, and after the model makes predictions, the data is restored and different-sourced belief degrees are fused, then loss calculation is performed with their corresponding labels and pseudo-labels to obtain corresponding losses $\mathcal{L}^{l_{2}}$ and $\mathcal{L}^{u}$ remaining consistent with the procedure in the pre-training stage.
@@ 2.3. Voxel-wise Asymptotic Learning
Inspired by curriculum learning (Bengio et al., 2009), we propose a strategy that utilizes voxel-level uncertainty to guide the model in gradually learning features. Initially, the model focuses on features that are easier to learn, asymptotically increasing the weight of voxels with higher uncertainty throughout the training process, thus enabling the model to pay further attention to features that are harder to learn. First, in the previous section, we obtained an updated uncertainty assessment for each voxel of the sample after fusion. Given the effectiveness of Shannon entropy in measuring probability distributions (Shannon, 1948; Wu et al., 2021), the uncertainty assessment for each voxel can further be specified as:
@@
where $\epsilon$ controls the amplitude of change of dynamic weight and $\zeta(h)=\frac{2h}{H}-1\in[-1,1],h=1,...,H$, $h$ represents the current epoch index and $H$ denotes the total number of training epochs. Moreover, $Sigmoid$ denotes the Sigmoid activation function, $\varsigma(s(z))=\frac{2s(z)}{Z}-1\in[-1,1],z=1,...,Z$, $s(z)$ represents the ordinal number of corresponding voxel $z$ which is obtained by sorting the uncertainty of all voxels owned by each sample of a different set of samples in a descending order. Therefore, for losses $\mathcal{L}^{l_{1}},\mathcal{L}^{l_{2}},\mathcal{L}^{u}$, we attach corresponding weights for fused predictions of each voxel on different samples, which can be given as:
(6) $\mathcal{L}^{\tau}_{w}=\sum_{k=1}^{G}\sum_{s(z)^{\tau}=1}^{Z}\varphi(h,s(z)^{\tau})\mathcal{L}^{\tau}_{k,s(z)}/Z$
where $\tau,k,G=(l_{1},i,A),(l_{2},i,A),(u,j,B)$, $\mathcal{L}_{k,s(z)}$ represents the loss of $z_{th}$ voxel in $k_{th}$ sample and $s(z)$ denotes corresponding ordinal number of voxel in $k_{th}$ sample. In sum, the total optimization objective $\mathcal{L}_{ini}$ and $\mathcal{L}_{s}$ for the initial network in pre-training stage and the student network in self-training stage can be given as:

### 11 Apr 2024 v2

@@ Abstract.
Figure 1. The diagram illustrates the fusion process of the model’s predictions for the voxels at the same location in the original samples and the samples after mixed area restoration. It demonstrates the redistribution and adjustment of the voxel’s confidence level in belonging to a category after interacting with the overall uncertainty of the prediction.
@@
  • The proposed method realizes the state-of-the-art (SOTA) performance on the task of semi-supervised medical image segmentation on four popular benchmark datasets and achieves performance that surpasses methods utilizing 20% labeled data with only 5% labeled data on TBAD dataset.
@@ 2.1. Generalized Probabilistic Framework
Evidential deep learning has achieved excellent results in multiple fields. We extend the framework of evidential deep learning to semi-supervised medical segmentation tasks by further improving the uncertainty measurement methods and probability assignments fusion rule of traditional evidence theory. We map the uncertainty measures of evidential deep learning to the multi-objective subsets of traditional evidence theory and fuse reliable information for both mixed and original sample annotation areas during the prediction phase, thereby achieving more robust model predictions with refined uncertainty measures. Moreover, drawing inspiration from curriculum learning (Bengio et al., 2009), we design a novel learning mechanism based on the refined uncertainty measure for the model to gradually focus on difficult-to-learn features.
In detail, let $g^{(x,y,z)}_{\mathcal{M}^{l}},g^{(x,y,z)}_{\mathcal{T}^{l_{1}}},g^{(x,y,z)}_{\mathcal{T}^{l_{2}}},g^{(x,y,z)}_{\mathcal{M}^{ul}},g^{(x,y,z)}_{\mathcal{T}^{u}}\in\mathcal{R}^{C}$ denote belief mass from evidential predictive results (Sensoy et al., 2018; Deng et al., 2023) from classifier of initial network $\mathcal{N}$ in pre-training stage on samples $\mathcal{M}^{l},\mathcal{T}^{l}$, and student network $\mathcal{N}_{s}$ in self-training stage on samples $\mathcal{T}^{l},\mathcal{M}^{ul},\mathcal{T}^{u}$ for voxel at position $(x,y,z)$ (sample indexes are omitted if there is no confusion). The final trained parameters of the initial network $\mathcal{N}$ will be utilized as the parameters of the teacher and student networks in the beginning of the self-training phase. The basic probability mass assignments for voxels can be defined as $\mathcal{P}^{(x,y,z)}=\{\{g^{(x,y,z)}(C_{n})\}_{n=0}^{N-1},u^{(x,y,z)}\}$ and $u^{(x,y,z)}=1-\sum_{n=0}^{N-1}g^{(x,y,z)}(C_{n})$ denotes the original uncertainty measure. The generalized form of predictive probability assignments can be formulated as:
@@ 2.2. Uncertainty-aware Evidential Fusion
In the previous section, we have mentioned that the method proposed in this article is divided into pre-training and self-training stages. For the pre-training stage, we first feed the samples into the initial network to obtain their predictive results. Due to the uniqueness of the mixed strategy (Bai et al., 2023a), we match the original labeled samples with the mixed samples one-to-one, and merge the evidential results of the corresponding areas of the two predictions. More specifically, if there are samples $E$ and $T$, which exchange features in the corresponding area $\mathcal{D}$ to obtain the mixed samples $E^{\prime}$ and $T^{\prime}$, after obtaining the evidential predictive results for the original samples $E$, $T$ and the mixed samples $E^{\prime}$, $T^{\prime}$, we return the exchanged areas of the mixed samples prediction back to their original positions and fuse them with the model’s predictive results for the original samples. Finally, the combined results and prediction of original data are used for loss calculation with the corresponding labels $Y^{l}$ to obtain $\mathcal{L}^{l_{1}}$, both of which consist of Dice loss and Cross-Entropy loss. We consider the fusion of voxel-level evidential results from two different predictive outputs that come from the same sample and are located at the same position $(x,y,z)$. For restored samples in $\mathcal{M}^{l}$ and original samples in $\mathcal{T}^{l}$, the proposed combination process for voxel at same position can be defined as:
@@
(3) $g_{fused}(C_{n})=g_{\mathcal{M}^{l}}(C_{n})g_{\mathcal{T}^{l}}(C_{n})+\frac{|C_{n}|}{|C_{N}|}\cdot$
$(g_{\mathcal{M}^{l}}(C_{n})g_{\mathcal{T}^{l}}(C_{N})+g_{\mathcal{T}^{l}}(C_{n})g_{\mathcal{M}^{l}}(C_{N}))$
where $n=0,...,N-1$, $|C_{n}|=1,|C_{N}|=N$ and $g_{fused}(C_{N})=g_{\mathcal{M}^{l}}(C_{N})*g_{\mathcal{T}^{l}}(C_{N})$. This fusion process evolves from the traditional Dempster’s combination rule requiring normalization (Dempster, 2008). The proposed fusion strategy maintains the design for the interaction between the confidence level of each category and the uncertain information, while simultaneously controlling the degree of interaction to avoid the confidence level of each category being dominated by uncertain information. Therefore, the interaction between the confidence associated with each category and the uncertain components is subject to certain restrictions, and we prefer the model to reduce uncertainty through the learning process, rather than directly converting uncertain parts into definite confidence assignments. If the fusion of probabilities follows the original Dempster’s combination rule directly, it may lead to the model being overly confident about certain predictions, thereby diminishing the focus on corresponding features, even though learning these features should still be encouraged. Similarly, for samples in the self-training stage, the fusion process they undergo is very similar to that in the pre-training stage. Labeled and unlabeled data undergo the same mixing operation as in the pre-training stage, and after the model makes predictions, the mixed samples are restored and different-sourced belief degrees are fused, then loss calculation is performed with their corresponding labels and pseudo-labels to obtain corresponding losses $\mathcal{L}^{l_{2}}$ and $\mathcal{L}^{u}$ remaining consistent with the procedure in the pre-training stage.
@@ 2.3. Voxel-wise Asymptotic Learning
Inspired by curriculum learning (Bengio et al., 2009; Zhang et al., 2024), we propose a strategy that utilizes voxel-level uncertainty to guide the model in gradually learning features. Initially, the model focuses on features that are easier to learn, asymptotically increasing the weight of voxels with higher uncertainty throughout the training process, thus enabling the model to pay further attention to features that are harder to learn. First, in the previous section, we obtained an updated uncertainty assessment for each voxel of the sample after fusion. Given the effectiveness of Shannon entropy in measuring probability distributions (Shannon, 1948; Wu et al., 2021), the uncertainty assessment for each voxel can further be specified as:
@@
where $\epsilon$ controls the amplitude of change of dynamic weight and $\zeta(h)=\frac{2h}{H}-1\in[-1,1],h=1,...,H$, $h$ represents the current epoch index and $H$ denotes the total number of training epochs. Moreover, $Sigmoid$ denotes the Sigmoid activation function, $\varsigma(s(z))=\frac{2s(z)}{Z}-1\in[-1,1],z=1,...,Z$, $s(z)$ represents the ordinal number of corresponding voxel $z$ which is obtained by sorting the uncertainty of all voxels owned by each sample in a descending order and $Z$ is the total number of voxels in a sample of a different groups of samples. Therefore, for losses $\mathcal{L}^{l_{1}},\mathcal{L}^{l_{2}},\mathcal{L}^{u}$, we attach corresponding weights for fused predictions of each voxel on different samples in diverse sample groups, which can be given as:
(6) $\mathcal{L}^{\tau}_{w}=\sum_{k=1}^{G}\sum_{z=1}^{Z}\varphi(h,s(z)^{\tau})\mathcal{L}^{\tau}_{k,z}/Z$
where $\tau,k,G=(l_{1},i,A),(l_{2},i,A),(u,j,B)$, $\mathcal{L}_{k,z}$ represents the loss of $z_{th}$ voxel in $k_{th}$ sample and $s(z)$ denotes corresponding ordinal number of voxel in $k_{th}$ sample ($Z$ changes when the group of samples are is different). In sum, the total optimization objective $\mathcal{L}_{ini}$ and $\mathcal{L}_{s}$ for the initial network in pre-training stage and the student network in self-training stage can be given as: